OpenStudy (anonymous):

What is the standard deviation of the following data set? 5, 10, 8, 6, 11 a. square root 26 b. square root 10 c. square root 5.2 d. square root 6.5

OpenStudy (amistre64):

i believe its: subtract the mean from the data entries square the results add them up and divide by the number of entries (-1 for a degree of freedom) then sqrt it

OpenStudy (amistre64):

\[var=\frac{\sum(x-\bar x)^2}{n-1}\]

OpenStudy (amistre64):

if this is a population then just divide by "n" if its a sample of a population; then we do the degree of freedom

OpenStudy (amistre64):

vqr = sd^2 sqrt(var)=sd

OpenStudy (amistre64):

typos abound :)

OpenStudy (anonymous):

so would it be c.?

OpenStudy (amistre64):

dunno

OpenStudy (anonymous):

amistre are you sure it's \( n-1 \) in the denominator?

OpenStudy (amistre64):

on this one, i think itll go with the population and ot the sample; so we can just "n" the denom

OpenStudy (anonymous):

im confused

OpenStudy (amistre64):

and NOT the sample ..... my keyboard hates me

OpenStudy (amistre64):

being confused is a rather vague and broad statement; how do we narrow it down?

OpenStudy (amistre64):

first, find the mean by adding up all the data points and dividning by how many there are; in this case we have 5?

OpenStudy (zarkon):

you need to determine if this is a sample or a population because the both answers are on the list

OpenStudy (amistre64):

5+10+8+6+11 = 40 40/6 = 6 2/3 right?

OpenStudy (anonymous):

if it helps, answer is d

OpenStudy (amistre64):

opps, 40/5 = 8 mean would be 8

OpenStudy (anonymous):

the answer is d ?

OpenStudy (anonymous):

@zarkon, i missed that. what does it mean?

OpenStudy (amistre64):

(5-8)^2 = 9 (10-8)^2 = 4 (8-8)^2 = 0 (6-8)^2 = 4 (11-8)^2 = 9 ------------ added: 26 26/5 = variance sqrt(variance) = sd

OpenStudy (amistre64):

assuming population

OpenStudy (zarkon):

the population std is c...the sample std is d

OpenStudy (amistre64):

sqrt(5.2) is what i end up with if i did the mathing right :)

OpenStudy (anonymous):

so is it d or c?

OpenStudy (amistre64):

it would help if you had something useful to contribute to the conversation ... imo

OpenStudy (amistre64):

the wording of the ? would assume population and not sample to me

OpenStudy (anonymous):

ah i see http://www.documentingexcellence.com/stat_tool/SD.htm learn something new every day! i would assume "population" here, but that is an assumption

OpenStudy (anonymous):

This seems to be the the standard deviation of the sample is the same as the population standard deviation so the denominator should be \(n \)

OpenStudy (anonymous):

See here: http://en.wikipedia.org/wiki/Standard_deviation#Estimation